# ObjectAdd: Adding Objects into Image via a Training-Free Diffusion Modification Fashion

### v2

@@ Abstract
We introduce ObjectAdd, a training-free diffusion modification method to add user-expected objects into user-specified area. The motive of ObjectAdd stems from: first, describing everything in one prompt can be difficult, and second, users often need to add objects into the generated image. To accommodate with real world, our ObjectAdd maintains accurate image consistency after adding objects with technical innovations in: (1) embedding-level concatenation to ensure correct text embedding coalesce; (2) object-driven layout control with latent and attention injection to ensure objects accessing user-specified area; (3) prompted image inpainting in an attention refocusing & object expansion fashion to ensure rest of the image stays the same. With a text-prompted image, our ObjectAdd allows users to specify a box and an object, and achieves: (1) adding object inside the box area; (2) exact content outside the box area; (3) flawless fusion between the two areas. Code is available at https://github.com/potato-kitty/ObjectAdd.
Keywords:
Diffusion Model Training-Free Text to Image Inpainting
1 Introduction
@@
One example is Midjourney, where users can create and share AI-generated images [17]. Another example is ChatGPT, of which the latest version allows users to request image generation [18]. Even classical software tools like Photoshop have improved their functionality by incorporating AI technology, enabling them to perform previously unimaginable tasks, such as changing the orientation of objects in photos [19]. These applications pave the way for the future of artificial general intelligence (AGI) as they can be operated solely through natural language commands. With the use of these tools, even individuals without any formal training can now create their own artwork, leading to a significant boost in the efficiency of the art industry. However, most text-to-image models rely solely on text prompts as inputs. This makes it challenging for users to convey spatial relationships between objects. Take the most famous large model, ChatGPT, for example. It requires users to describe all their requirements through chatting, which is not an easy task for non-professional personnel. Some approaches have been proposed to solve this problem, with ControlNet [13] possibly being one of the most famous models. ControlNet allows users to include canny edge or human pose as extra inputs, enabling control over the pose of humans or animals in the generated image. With the use of ControlNet, users can control the generated result in a more intuitive way, but it still requires users to decide the layout all at once before seeing any result.
@@
Thus, we conclude two practical issues: first, describing everything in a single prompt can be difficult, and second, users often need to modify the generated image multiple times to achieve their desired results. Unfortunately, current models are stuck in these challenges. For example, as shown in Fig. 1, in our tests with ChatGPT-4, which uses DALL-E 3 for image generation, we find that while the initial images are of high quality, they lose consistency after modification. Details such as the orientation of a person and their clothing are lost.
@@
2 Related Works
2.1 Training-Free Control
@@
2.2 Training-Dependent Image Editing
Though fine-tuning a large model could be difficult, an external model can be trained to assist the larger model. For instance, in ITI-Gen [10], researchers train an encoder to obtain distinctive token embeddings. This is done to accurately generate human faces that match the descriptive words in the prompt, such as skin tone. Another example is seen in GLIGEN [11], where an extra self-attention layer is trained and integrated into the pretrained model. This layer encodes and transmits location information to the model. The benefit of this method is its stability compared to training-free approaches. However, it requires a dataset tailored to its specific task, which can sometimes be challenging to obtain.
2.3 Decoupled Text Embedding
@@
3 Methodology
3.1 Preliminaries
@@
Herein, $H$, $W$ and $C$ respectively indicate image height, width and channel number. Inside the stable diffusion model $F$, there are some intermediate representations that are the focus in our method: (1) Text embedding $\mathcal{E}_{\mathcal{P}}\in\mathbb{R}^{N\times D}$, usually derived from a CLIP model [1], where $N$ represents the maximum token sequence length in the text encoder and $D$ denotes the embedding dimension; (2) The $t$-th step latent denoised representation $\mathcal{I}^{t}\in\mathbb{R}^{H\times W\times C}$; (3) The $\gamma$-layer cross-attention map at the $t$-th step $\mathcal{C}^{\gamma t}\in\mathbb{R}^{N\times(H_{\gamma}\cdot W_{\gamma})}$, with $\mathcal{C}^{\gamma t}_{i,:}\in\mathbb{R}^{H_{\gamma}\cdot W_{\gamma}}$ indicating the attentive score of the $\gamma$-th layer output to the $i$-th token. For ease of the following representation, we reshape $C^{{\gamma}t}_{i,:}$ into $H_{\gamma}\times W_{\gamma}$. Notice these symbols will be revised accordingly in the following sections, in order to adapt to outputs of different prompts. For example, $\mathcal{C}^{{\gamma}t}$ will be revised as $\bar{\mathcal{C}}^{{\gamma}t}$ for adapting to Eq. (2) and $\tilde{\mathcal{C}}^{{\gamma}t}$ for adapting to Eq. (4), and many others.
Our intention in this paper, is to edit on image $\mathcal{I}^{0}$ to derive a user-expected version ¯$\bar{\mathcal{I}}^{0}$ under two application-oriented settings: 1) a user-drawn box in image $\mathcal{I}^{0}$ and 2) a user-specified word-level prompt $\mathcal{W}$ describing an object. The box can be characterised as a binary mask $\mathcal{M}\in\{0,1\}^{H\times W}$ where $\mathcal{M}_{i,j}=1$ if pixel $\mathcal{I}^{0}_{i,j}$ falls into the drawn box, and $\mathcal{M}_{i,j}=0$ else. Note that, the $\mathcal{M}$ is downsampled into $\mathcal{M}^{\gamma}\in\mathbb{R}^{H_{\gamma}\times W_{\gamma}}$, for adapting to the $\gamma$-th layer image shape.
@@
3.2 Text Embedding Coalesce
We first deal with the coalesce of prompts $\mathcal{P}$ and $\mathcal{W}$ in Eq. (2). As specified in Sec. 2.3, the self-attention mechanism in text encoder causes embedding confusion if simply concatenating two prompts together $concat(\mathcal{P},\mathcal{W})$. Taking $\mathcal{P}$ = “A woman wearing glasses” and $\mathcal{W}$ = “A hat” as a simple example, the concatenation results in prompt “A woman wearing glasses A hat” tokenized as (“[CLS],” “a,” “woman,” “wearing,” “glasses,” “a,” “hat,” “[SEP]”) where “[CLS]” and “[SEP]” respectively denote the start and end tokens. For self-attention without causal masks, these tokens are mutually attentioned therefore causing different embeddings of “a,” “woman,” “wearing,” “glasses,” “a,” and “hat,” compared to individually embed prompts $\mathcal{P}$ (“[CLS],” “a,” “woman,” “wearing,” “glasses,” “[SEP]”) and $\mathcal{W}$ (“[CLS],” “a,” “hat,” “[SEP]”). Similarly, masked self-attention causes different embeddings of “a” and “hat.” Most current diffusion models deploy a masking-free CLIP-style text encoder [1], which contradicts the intention in this paper that users add an object into a selected area of an image, while maintaining the rest content almost unchanged.
Instead of having an operation on the inputs of the text encoder, to reach correct text embedding coalesce, we choose to modify outputs of the text encoder, i.e., text embeddings. We embed prompt $\mathcal{P}$ and prompt $\mathcal{W}$ separately, as $\mathcal{E}_{\mathcal{P}}\in\mathbb{R}^{N\times D}$ and $\mathcal{E}_{\mathcal{W}}\in\mathbb{R}^{N\times D}$. Recall $N$ (usually $N=77$ in CLIP text encoder) denotes the maximum token sequence length in the text encoder. Denoting the actual token (no “[PAD]”, “[CLS]” and “[EOS]”) number of prompt $\mathcal{P}$ as $N_{\mathcal{P}}$, and $N_{\mathcal{W}}$ for prompt $\mathcal{W}$, we formulate the text embedding coalesce as:
@@ 3.2 Text Embedding Coalesce
3.3 Object-Driven Layout Control
@@
3.3.1 Latent Information Injection.
To inject latent-level information, we additionally perform an individual denoising process for the word-level prompt $\mathcal{W}$:
@@
3.3.2 Attention Information Injection.
@@
Note that, we apply information injection across all layers, but only in the first few time steps, to avoid inconsistent style, abrupt edge, etc. Specifically, as $t=T\rightarrow 0$, latent information injection is applied in the first 20% time steps while attention information injection is applied in the first 30% time steps. Details will be ablated in the Appendix.
3.4 Prompted Image Inpainting
Utilizing text embedding coalescence and object-driven layout control, we add a high-quality object to the original image. A prime example is illustrated between Fig. 2 and Fig. 2, where the prompt $\mathcal{W}$ = “A hat” is seamlessly integrated into the user-drawn area. However, Fig. 2 also serves as a negative example, displaying inconsistencies in background, face, and gesture. To maintain consistency in the out-of-box area, in Fig. 2, we employ attention refocusing and object expansion to shift the mask attention from user-drawn box to generated object.
3.4.1 Attention Refocusing.
Our motive mainly originates from an in-depth analysis of the cross-attention map $\bar{\mathcal{C}}_{k,:}^{\gamma t}$ for the added object as visualized in Fig. 3. In the middle diffusion steps typically $0.3\cdot T<t<0.5\cdot T$, we find the central area of $\mathcal{W}$ described object will be well focused in the layer of $(H_{\gamma},W_{\gamma})=(32,32)$. However, the edge area would be mixed up with the out-of-box area. Therefore, we consider refocusing the attention pixel, to adjust a user-drawn box, i.e., the mask $\mathcal{M}^{\gamma}$, into the object-centralized area ¯$\bar{\mathcal{M}}^{\gamma}$.
To this end, we perform clustering upon the cross-attention map $\bar{\mathcal{C}}_{k,:}^{\gamma t}$ to segment it into $K$ sub-figures, as marked with different colors in Fig. 3:
@@ 3.4 Prompted Image Inpainting
where $\mathcal{H}_{1}$ is a threshold. Fig. 3 manifests the object-centralized area.
Finally, in Fig. 3, we refocus the user-drawn area by updating the binary mask as ¯$\bar{\mathcal{M}}^{\gamma}$ where ¯$\bar{\mathcal{M}}^{\gamma}_{i,j}=1$ if attention pixel ($i,j$) falls into object-centralized area and 0 otherwise. For better performance, in Fig. 3, we apply mathematical morphology [15] to remove small objects and fill holes inside the masked area.
3.4.2 Object Expansion.
@@
$\begin{split}D(\bar{\mathcal{I}}_{x,y}|\mathcal{I}_{i,j},\mathcal{S}_{\mathcal{I}_{i,j}})&=0.5\cdot dist(\bar{\mathcal{I}}_{x,y},\;\bar{\mathcal{I}}_{i,j})+0.5\cdot dist\big{(}\bar{\mathcal{I}}_{x,y},\;avg(\mathcal{S}_{\mathcal{I}_{i,j}})\big{)},\\ &\qquad\;s.t.\quad\bar{\mathcal{M}}_{i,j}==1,\;\bar{\mathcal{M}}_{x,y}==0,\\ &x\in\{i,i-1,i+1\},\;y\in\{y,y-1,y+1\},\end{split}$ (13)
@@
After this step, the ¯$\bar{\mathcal{I}}^{t}$ is now a combination of detected object and the non-object area of original prompt derived latent $\mathcal{I}^{t}$. We apply the above swapping at a random time step between $0.3\cdot T<t<0.5\cdot T$. Since we apply this operation in a time step far away from diffusion end, the robustness of diffusion model would make up the missed detected area like the smaller area in Fig. 3 and unnatural edges. Details will be ablated in the Appendix.
4 Experimentation
4.1 Experimental Setup
4.1.1 Implementation Details.
We conduct experiments on two Nvidia GeForce RTX 3090 GPUs. In the case of object-driven layout control, as the diffusion step $t=T\rightarrow 0$, we incorporate latent information during the initial 20% of time steps, while the injection of attention information occurs within the first 30% of time steps. We employ the strategies of attention refocusing and object expansion to recover the prompted image at the step of $t=15$, adjustable provided $0.3\cdot T<t<0.5\cdot T$. Pertaining to the clustering operation in attention refocusing, we establish the cluster number $K=6$ in Eq. (9) and set the threshold $\mathcal{H}_{1}=0.35$ in Eq. (11) along with the parameter $\mathcal{H}_{2}=5$ in Eq. (14).
4.1.2 Compared Methods.

### v5

@@ Abstract
We introduce ObjectAdd, a training-free diffusion modification method to add user-expected objects into user-specified area. The motive of ObjectAdd stems from: first, describing everything in one prompt can be difficult, and second, users often need to add objects into the generated image. To accommodate with real world, our ObjectAdd maintains accurate image consistency after adding objects with technical innovations in: (1) embedding-level concatenation to ensure correct text embedding coalesce; (2) object-driven layout control with latent and attention injection to ensure objects accessing user-specified area; (3) prompted image inpainting in an attention refocusing & object expansion fashion to ensure rest of the image stays the same. With a text-prompted image, our ObjectAdd allows users to specify a box and an object, and achieves: (1) adding object inside the box area; (2) exact content outside the box area; (3) flawless fusion between the two areas. The project is at https://github.com/potato-kitty/ObjectAdd.
Index Terms:
Diffusion Model, Training-Free, Text to Image, Inpainting
I Introduction
@@
One example is Midjourney, where users can create and share AI-generated images [6]. Another example is ChatGPT, of which the latest version allows users to request image generation [7]. Even classical software tools like Photoshop have improved their functionality by incorporating AI technology, enabling them to perform previously unimaginable tasks, such as changing the orientation of objects in photos [8]. These applications pave the way for the future of artificial general intelligence (AGI) as they can be operated solely through natural language commands. With the use of these tools, even individuals without any formal training can now create their own artwork, leading to a significant boost in the efficiency of the art industry. However, most text-to-image models rely solely on text prompts as inputs. This makes it challenging for users to convey spatial relationships between objects. Take the most famous large model, ChatGPT, for example. It requires users to describe all their requirements through chatting, which is not an easy task for non-professional personnel. Some approaches have been proposed to solve this problem, with ControlNet [9] possibly being one of the most famous models. ControlNet allows users to include canny edge or human pose as extra inputs, enabling control over the pose of humans or animals in the generated image. With the use of ControlNet, users can control the generated result in a more intuitive way, but it still requires users to decide the layout all at once before seeing any outcome result.
Thus, we conclude two practical issues: first, describing everything in a single prompt can be difficult, and second, users often need to modify the generated image multiple times to achieve their desired results. Unfortunately, current models are stuck in these challenges. For example, as shown in Fig. 1, in our tests with ChatGPT-4, which uses DALL-E 3 for image generation, we find that while the initial images are of high quality, but they lose consistency after modification. Details such as the orientation of a person and their clothing are lost.
@@
For user convenience, we also provide an extension of our method that allows real images as the input. To achieve this, we first reshape the input real image to fit the user-selected area and then apply a newly proposed inversion method [] before utilizing our ObjectAdd methods. While all other modifications remain the same as in the case of non-real image inputs, we replace the mask for the user-selected area with the segmentation of the reshaped input real image. Our experiments indicate that this approach results in improved performance.
II Related Works
II-A Training-Free Control
@@
II-B Training-Dependent Image Editing
Though fine-tuning a large model could be difficult, an external model can be trained to assist the larger model. For instance, in ITI-Gen [18], researchers train an encoder to obtain distinctive token embeddings. This is done to accurately generate human faces that match the descriptive words in the prompt, such as skin tone. Another example is seen in GLIGEN [19], where an extra self-attention layer is trained and integrated into the pretrained model. This layer encodes and transmits location information to the model. The benefit of this method is its stability compared to training-free approaches. However, it requires a dataset tailored to its specific task, which can sometimes be challenging to obtain. InstructPix2Pix [20] could be the one that has a most similar goal to our approach. It allows users to specify their desired modifications to an input image through prompts. This model takes input image as conditional vector, as well as the instruction prompt indicated by user. The output of this approach is the edit version of input image, following the instruction. However, it still only allows prompts as input for specifying the requirements, and its results often fail to guarantee the consistency.
II-C Decoupled Text Embedding
@@
II-D Diffusion Based Real Image Editing
Diffusion models have demonstrated significant capabilities in controllable image generation [12, 22]. There is a growing interest in using them for real image editing. Since diffusion models start with noise as input, the primary challenge to edit real images is to convert the input real image into noise in such a way that the noise can be reconstructed into the original image after the denoising steps. This conversion process is called inversion. The most commonly used method today is DDIM inversion [23], which inverts each denoising step by adding the predicted noise to the latent representation. Several approaches have been proposed to improve the quality of inversion. In [24], the researchers suggested that inversion could be considered a fixed-point iteration problem and could be solved using iterative updating methods. ReNoise [25] addresses the inversion problem from a similar perspective, but introduces minor biases in each step to enhance the edit-ability of the inversion results.
Once the real image has been inverted to noise, it can be used as the input for the diffusion model. Various methods can then be applied to control the generation process, thereby editing the real image. MasaCtrl [22] is one example. In that paper, researchers replace information in the editing branch with the corresponding information in the reconstruction branch to maintain consistency while using new prompts as conditions for generation. This approach allows editing original image, applying modifications while retaining most of the original information.
III Methodology
III-A Preliminaries
@@
Herein, $H$, $W$ and $C$ respectively indicate image height, width and channel number. Inside the stable diffusion model $F$, there are some intermediate representations that are the focus in our method: (1) Text embedding $\mathcal{E}_{\mathcal{P}}\in\mathbb{R}^{N\times D}$, usually derived from a CLIP model [2], where $N$ represents the maximum token sequence length in the text encoder and $D$ denotes the embedding dimension; (2) The $t$-th step latent denoised representation $\mathcal{I}^{t}\in\mathbb{R}^{H\times W\times C}$; (3) The $\gamma$-layer cross-attention map at the $t$-th step $\mathcal{C}^{\gamma t}\in\mathbb{R}^{N\times(H_{\gamma}\cdot W_{\gamma})}$, with $\mathcal{C}^{\gamma t}_{i,:}\in\mathbb{R}^{H_{\gamma}\cdot W_{\gamma}}$ indicating the attentive score of the $\gamma$-th layer output to the $i$-th token. For ease of the following representation, we reshape $C^{{\gamma}t}_{i,:}$ into $H_{\gamma}\times W_{\gamma}$. Notice these symbols will be revised accordingly in the following sections, in order to adapt to outputs of different prompts. For example, $\mathcal{C}^{{\gamma}t}$ will be revised as $\bar{\mathcal{C}}^{{\gamma}t}$ for adapting to Eq. (2) and $\tilde{\mathcal{C}}^{{\gamma}t}$ for adapting to Eq. (4).
Our intend to edit on image $\mathcal{I}^{0}$ to derive a user-expected version ¯$\bar{\mathcal{I}}^{0}$ under two application-oriented settings: 1) a user-drawn box in image $\mathcal{I}^{0}$ and 2) a user-specified word-level prompt $\mathcal{W}$ describing an object. The box can be characterised as a binary mask $\mathcal{M}\in\{0,1\}^{H\times W}$ where $\mathcal{M}_{i,j}=1$ if pixel $\mathcal{I}^{0}_{i,j}$ falls into the drawn box, and $\mathcal{M}_{i,j}=0$ else. Note that, the $\mathcal{M}$ is downsampled into $\mathcal{M}^{\gamma}\in\mathbb{R}^{H_{\gamma}\times W_{\gamma}}$, for adapting to the $\gamma$-th layer image shape.
@@
III-B Text Embedding Coalesce
We first deal with the coalesce of prompts $\mathcal{P}$ and $\mathcal{W}$ in Eq. (2). As specified in Sec. II-C, the self-attention mechanism in text encoder causes embedding confusion if simply concatenating two prompts together $concat(\mathcal{P},\mathcal{W})$. Taking $\mathcal{P}$ = “A woman wearing glasses” and $\mathcal{W}$ = “A hat” as a simple example, the concatenation results in prompt “A woman wearing glasses A hat” tokenized as (“[CLS],” “a,” “woman,” “wearing,” “glasses,” “a,” “hat,” “[SEP]”) where “[CLS]” and “[SEP]” respectively denote the start and end tokens. For self-attention without causal masks, these tokens are mutually attentioned therefore causing different embeddings of “a,” “woman,” “wearing,” “glasses,” “a,” and “hat,” compared to individually embed prompts $\mathcal{P}$ (“[CLS],” “a,” “woman,” “wearing,” “glasses,” “[SEP]”) and $\mathcal{W}$ (“[CLS],” “a,” “hat,” “[SEP]”). Similarly, masked self-attention causes different embeddings of “a” and “hat.” Most current diffusion models deploy a masking-free CLIP-style text encoder [2], which contradicts the intention in this paper that users add an object into a selected area of an image, while maintaining the rest content unchanged.
Instead of operating the inputs of the text encoder, to reach correct text embedding coalesce, we modify outputs of the text encoder, i.e., text embeddings. We embed prompt $\mathcal{P}$ and prompt $\mathcal{W}$ separately, as $\mathcal{E}_{\mathcal{P}}\in\mathbb{R}^{N\times D}$ and $\mathcal{E}_{\mathcal{W}}\in\mathbb{R}^{N\times D}$. Recall $N$ (usually $N=77$ in CLIP text encoder) denotes the maximum token sequence length in the text encoder. Denoting the actual token (no “[PAD]”, “[CLS]” and “[EOS]”) number of prompt $\mathcal{P}$ as $N_{\mathcal{P}}$, and $N_{\mathcal{W}}$ for prompt $\mathcal{W}$, we formulate the text embedding coalesce as:
@@ III-B Text Embedding Coalesce
III-C Object-Driven Layout Control
@@
III-C1 Latent Information Injection
To inject latent-level information, we perform an individual denoising process for the word-level prompt $\mathcal{W}$:
@@ III-C1 Latent Information Injection
III-C2 Attention Information Injection
@@
We apply information injection across all layers, but only in the first few time steps, to avoid inconsistent style, abrupt edge, etc. Specifically, as $t=T\rightarrow 0$, latent information injection is applied in the first 20% time steps while attention information injection is applied in the first 30% time steps.
III-D Prompted Image Inpainting
Utilizing text embedding coalescence and object-driven layout control, we add a high-quality object to the original image. A prime example is illustrated between Fig. 2a and Fig. 2b, where the prompt $\mathcal{W}$ = “A hat” is seamlessly integrated into the user-drawn area. However, Fig. 2b also serves as a negative example, displaying inconsistencies in background, face, and gesture. To maintain consistency in the out-of-box area, in Fig. 2c, we employ attention refocusing and object expansion to shift the mask attention from user-drawn box to generated object.
III-D1 Attention Refocusing
Our motive mainly originates from an in-depth analysis of the cross-attention map $\bar{\mathcal{C}}_{k,:}^{\gamma t}$ for the added object as visualized in Fig. 3a. In the middle diffusion steps typically $0.3\cdot T<t<0.5\cdot T$, we find the central area of $\mathcal{W}$ described object will be well focused in the layer of $(H_{\gamma},W_{\gamma})=(32,32)$. However, the edge area would be mixed up with the out-of-box area. Therefore, we consider refocusing the attention pixel, to adjust a user-drawn box, i.e., the mask $\mathcal{M}^{\gamma}$, into the object-centralized area ¯$\bar{\mathcal{M}}^{\gamma}$.
To this end, we perform clustering upon the cross-attention map $\bar{\mathcal{C}}_{k,:}^{\gamma t}$ to segment it into $K$ sub-figures, as marked with different colors in Fig. 3b:
@@ III-D1 Attention Refocusing
where $\mathcal{H}_{1}$ is a threshold. Fig. 3c manifests the object-centralized area.
Finally, in Fig. 3d, we refocus the user-drawn area by updating the binary mask as ¯$\bar{\mathcal{M}}^{\gamma}$ where ¯$\bar{\mathcal{M}}^{\gamma}_{i,j}=1$ if attention pixel ($i,j$) falls into object-centralized area and 0 otherwise. For better performance, in Fig. 3e, we apply mathematical morphology [26] to remove small objects and fill holes inside the masked area.
III-D2 Object Expansion
@@
$\begin{split}D(\bar{\mathcal{I}}_{x,y}|\mathcal{I}_{i,j},\mathcal{S}_{\mathcal{I}_{i,j}})=\\ 0.5(\cdot dist(\bar{\mathcal{I}}_{x,y},\;\bar{\mathcal{I}}_{i,j})+dist\big{(}\bar{\mathcal{I}}_{x,y},\;avg(\mathcal{S}_{\mathcal{I}_{i,j}})\big{)}),\\ \qquad\;s.t.\quad\bar{\mathcal{M}}_{i,j}==1,\;\bar{\mathcal{M}}_{x,y}==0,\\ x\in\{i,i-1,i+1\},\;y\in\{y,y-1,y+1\},\end{split}$ (13)
@@
After this step, the ¯$\bar{\mathcal{I}}^{t}$ is now a combination of detected object and the non-object area of original prompt derived latent $\mathcal{I}^{t}$. We apply the above swapping at a random time step between $0.3\cdot T<t<0.5\cdot T$. Since we apply this operation in a time step far away from diffusion end, the robustness of diffusion model would make up the missed detected area like the smaller area in Fig. 3f and unnatural edges. Details will be ablated in Sec. IV-D.
Figure 4: Pre-processing of input real image.
III-E Real Image Editing
To apply ObjectAdd on real images, we pre-process this extra input, using the results to update the mask and generate the new input in Eq. (4). We omit the layout control methods discussed in Sec. III-C since we can directly resize the input.
For the pre-processing of the input real image, as shown in Fig. 4, we first segment the object and then resize the segmentation to the user-chosen area, as illustrated from Fig. 4a to Fig. 4b. We represent the results of this step as $\tilde{\mathcal{Z}}_{0}$. For convenience, we choose pictures with a white background as examples, allowing for simple segmentation algorithms. For input images with complex backgrounds, a pre-trained segmentation model might be required. We then apply inversion towards $\tilde{\mathcal{Z}}_{0}$ to obtain the noise $\tilde{\mathcal{Z}}_{T}$. We use $\tilde{\mathcal{Z}}_{T}$ to replace the random noise $\mathcal{I}_{T}$ in Eq. (4), resulting in the following updated equation:
~$\tilde{\mathcal{I}}^{0}=F(\tilde{\mathcal{Z}}_{T};\mathcal{W}).$ (16)
In this way, we ensure that the object in ~$\tilde{\mathcal{I}}^{0}$ is generated in the user-specified location after denoising, fulfilling the role of layout control methods discussed in Sec. III-C.
Another modification for real image editing is that, since we have the segmentation of the input, we can generate a more precise mask from it, as shown from Fig. 4b to Fig. 4c. We denote this new mask as $\mathcal{M}^{\prime}$. Compared to the original mask $\mathcal{M}$, which is represented by the red rectangle area in Fig. 3a, $\mathcal{M}^{\prime}$ is more accurate. Thus, in real image editing tasks, the mask $\mathcal{M}$ mentioned in previous subsections will be replaced with $\mathcal{M}^{\prime}$.
For using whatever inversion methods, there are always errors during both inverse and denosing process. For this reason, we inject the corresponding inverted latent $\tilde{\mathcal{Z}}_{\tilde{t}}$ at time-step $\tilde{t}$ where $\tilde{t}$ is set to 39 in this paper, the detail of such injection is as following:
¯$\bar{\mathcal{I}}^{\tilde{t}}=(1-\mathcal{M}^{\prime})\odot\bar{\mathcal{I}}^{\tilde{t}}+\mathcal{M}^{\prime}\odot\tilde{\mathcal{Z}}_{\tilde{t}}.$ (17)
IV Experimentation
IV-A Experimental Setup
IV-A1 Implementation Details
We conduct experiments on one Nvidia GeForce RTX 3090 GPU. In the case of object-driven layout control, as the diffusion step $t=T\rightarrow 0$, we incorporate latent information during the initial 20% of time steps, while the injection of attention information occurs within the first 30% of time steps. We employ the strategies of attention refocusing and object expansion to recover the prompted image at the step of $t=15$, adjustable provided $0.3\cdot T<t<0.5\cdot T$. Pertaining to the clustering operation in attention refocusing, we establish the cluster number $K=6$ in Eq. (9) and set the threshold $\mathcal{H}_{1}=0.35$ in Eq. (11) along with the parameter $\mathcal{H}_{2}=5$ in Eq. (14).
IV-A2 Compared Methods